OpenStudy (anonymous):

because 1/3+1 it is = 4/3 ?????????

OpenStudy (anonymous):

yes thats correct think of it as 3/3 + 1/3 = 4/3

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 1+\frac{1}{3} }\)\(\Huge\color{blue}{ \sf = \frac{3}{3}+\frac{1}{3} }\)\(\Huge\color{blue}{ \sf =\frac{3+1}{3} }\)\(\Huge\color{blue}{ \sf = \frac{4}{3} }\)

OpenStudy (solomonzelman):

or you can just put them together (literally put them together)

OpenStudy (anonymous):

Now I get it ... thanks guys

OpenStudy (solomonzelman):

No problem :)